x^2+2x=3 Solve by graphing
Found 2 solutions by Alan3354, jorel1380:
Answer by Alan3354(69443)

dl the FREE graph software at
http://www.padowan.dk
--------------
x^2+2x=3
x^2 + 2x - 3
Use Insert

Answer by jorel1380(3719)

x^2+2x=3
x2+2x-3=0
(x+3)(x-1)=0
x=-3 or 1